OpenStudy (calculator):

What mass of solid residue can be obtained from the thermal decomposition of 4.10 g of anhydrous calcium nitrate? Ans: 1.40g

OpenStudy (calculator):

The formula should be Ca(NO3)2-H2O Mr of Ca(NO3)2 is 164g/mol Total Mr is 182 g/mol 164/182 x 4.1 = 3.69 g I can't get 1.4 g

OpenStudy (calculator):

@ghazi

OpenStudy (jfraser):

probably because you're not driving off water, since it's already anhydrous. You're decomposing the substance, probably into calcium oxide and NO2 gas

OpenStudy (calculator):

@JFraser so you mean the residue should be calcium oxide.

OpenStudy (jfraser):

that would be my guess

OpenStudy (calculator):

Hmm but how do you know that Ca(NO3)2-H2O will decompose to calcium oxide and NO2 gas? its kinda hard to think of the products formed for me.

OpenStudy (jfraser):

take the water out of that formula, the problem specifically says "anhydrous", so there's no water

OpenStudy (calculator):

I get \[\frac{56}{182} \times 4.1=1.26g\]

OpenStudy (jfraser):

In predicting the product, I'm using the experience that calcium carbonate, CaCO3, decomposes into CaO and CO2, and guessing the rest. Calcium nitrate decomposing probably looks like:\[Ca(NO_3)_2 \rightarrow CaO(s) + N_xO_y\] the exact formula of the nitrogen gas isn't important, because the ratio of calcium nitrate to calcium oxide will be 1:1

OpenStudy (calculator):

oh i thought anhydrous is the one with water

OpenStudy (jfraser):

putting the prefix "an-" in front of a word negates its meaning, so an anhydrous salt is one with no water.

OpenStudy (calculator):

k got it, so about the calculation...

OpenStudy (jfraser):

you're using the wrong total mass, it should be \[\frac{56g CaO}{164g Ca(NO_3)_2} * 4.1g Ca(NO_3)_2 = 1.4g\]

OpenStudy (calculator):

got it already 1.4g,

OpenStudy (calculator):

yea thanks!